Using mold concept, find the concentration of NaOH using 0.10 moldm2 given that the volume of acid used was 12.0cm volume of Base 25cm2


1
Expert's answer
2021-08-04T07:13:43-0400

Using mole concept

Given that the chemical reaction is;

"NaOH +HCl\to NaCl + H_2 O"

"C_A = 0.10 \textsf{ mol/dm³}\\
V_A = 12.0cm³\\
V_B = 25cm³\\
C_B = ?"


"\dfrac{C_AV_A}{C_BV_B} = \dfrac{n_A}{n_B}"


"\dfrac{0.10 × 12}{C_B × 25} = \dfrac11"


"C_B = \dfrac{0.10× 12}{25} = 0.048 \textsf{ mol/dm³}"